# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ Abstract
We consider sequences of finite weighted random graphs that converge locally to unimodular i.i.d. weighted random trees. When the weights are atomless, we prove that the matchings of maximal weight converge locally to a matching on the limiting tree. For this purpose, we introduce and study unimodular matchings on weighted unimodular random trees as well as a notion of optimality for these objects. In this context, we prove that, in law, there is a unique optimal unimodular matching for a given unimodular tree. We then prove that this law is the local limit of the sequence of matchings of maximal weight. Along the way, we also show that this law is characterised by an equation derived from a message passing algorithm.
@@ 1 Introduction
Optimal matchings in random graphs is a problem with a long history dating back to the paper by Karp and Sipser [4568355]. In this seminal paper, they obtain the asymptotic size of a maximal matching in a sparse Erdős-Rényi random graph with average degree $c\in(0,+\infty)$. Their method relies on the analysis of a greedy algorithm that constructs a matching in a graph. It turns out that the algorithm is very close to optimality when $c\leq e$, and less so when $c>e$ (even if it is still asymptotically optimal), making the analysis trickier in this regime.
@@
More recently, Gamarnik, Nowick and Swirscsz [gamarnik2003maximum] studied the edge-weighted version of this maximal matching problem on the Erdős-Rényi random graph and on regular random graphs. Their approach does not rely on the Karp–Sipser algorithm, but on a local optimality ansatz inspired by the earlier introduction of the so-called objective method by Aldous [Aldous2004, aldous2018processes, Aldous1992AsymptoticsIT, aldous2000zeta2]. Gamarnik, Nowick and Swirscsz were able to compute the asymptotic weight of the optimal matching in Erdős-Rényi random graph when $c\leq e$ and in the special case where the weights have exponential distribution. These restrictions come from the fact that their method depends on the uniqueness of the solution of a fixed point in a belief propagation equation.
@@
Strikingly we realised that, under mild assumptions, this uniqueness is not required for the convergence in law of the optimal matching on a locally tree-like edge-weighted random graph. Informally, the limit $\mathbb{M}$ is a random matching on the unimodular limiting tree $(\mathbb{T},o)$ with iid edge-weights $w=(w(e))_{e\in E(\mathbb{T})}$ with law $\omega$, whose law is invariant by re-rooting. We call such random matchings unimodular (see Section 2 for precise definitions). In addition, the limit $\mathbb{M}$ is optimal in the sense that the quantity
@@
$\operatorname{perf}(\mathbb{T},o,w,\mathbb{M})$ $:=\mathbb{E}\left[\sum_{v\sim o}w(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}}\right]$ (1)
is maximal among all unimodular matchings of $(\mathbb{T},o,w)$.
@@
Before stating our main result, we need to introduce Unimodular Bienaymé Galton Watson random trees (UBGW), which are the limiting unimodular random trees of our graphs. Let $\pi$ be a distribution on non-negative integers with finite variance and generating function $\phi$. We say that a rooted tree $(\mathbb{T},{o})$ is a UBGW with offspring distribution $\pi$ if it has the following law:
  • The number of children of vertices of $\mathbb{T}$ are all independent.
  • The number of children of the root $o$ is distributed according to $\pi$.
  • The number of children of every non-root vertex is distributed according to $\hat{\pi}$, the sized biased version of $\pi$. The generating function of $\hat{\pi}$ is $\hat{\phi}(x)=\frac{\phi^{\prime}(x)}{\phi^{\prime}(1)}$.
We can now state our main result:
Theorem 1.
@@
Let $G_{n}=(V_{n},E_{n},w_{n})$ be a sequence of weighted random graphs and let $\mathbb{M}_{\mathrm{opt}}(G_{n})$ be a maximum matching on $G_{n}$. Choose a root ${o}_{n}$ uniformly among the vertices and suppose that the sequence of rooted weighted random graphs $(G_{n},{o}_{n})$ converges locally to a vertex-rooted Unimodular Bienaymé-Galton-Watson (UBGW) tree $(\mathbb{T},{o})$ with reproduction law $\pi$ and i.i.d edge-weights with law $\omega$ such that:
@@
Then there exists a unique (in law) optimal matching $\mathbb{M}$ on $(\mathbb{T},{o})$ such that $(G_{n},o_{n},\mathbb{M}_{\mathrm{opt}}(G_{n}))$ converges locally to $(\mathbb{T},{o},\mathbb{M})$. Furthermore the distribution of $(\mathbb{T},{o},\mathbb{M})$ is explicitly described in Proposition 9 (ii).
Remark.
As often with unimodular local limits, we also get a similar statement when the graphs are rooted at a uniform oriented edge. See Section 2 for precise statements.
The description of $\mathbb{M}$ given in Proposition 9 (ii) relies on the belief propagation mentioned above. In the setting of finite trees, this procedure always constructs an optimal matching. A good portion of this work, done in Section 3 is then to show that the same procedure produces an optimal random matching when it is applied to a weighted UBGW tree. The second important step, done in Section 4, is to show that it is possible to exhibit a matching of the (large) finite graph whose average weight per vertex is close to the performance of the limit as defined in (1).
To gain more insight into the belief propagation procedure and our results, we introduce informally the belief variables $Z(u,v)$ associated with oriented edges $(u,v)$ of a finite deterministic tree $T$. This will be done in detail in Section 3.1. Each variable $Z(u,v)$ captures the gain of allowing $v$ to be matched in $T_{v}$, the connected component of $T\setminus\{\{u,v\}\}$ containing $v$. When $T_{v}$ is finite, this corresponds to the difference between the maximal weight of a matching of $T_{v}$ and the maximal weight of a matching of $T_{v}\setminus\{v\}$. From this, we can deduce the so-called decision rule of the belief propagation procedure:
$\{u,v\}\in\mathbb{M}\Leftrightarrow Z(u,v)+Z(v,u)<w(u,v).$ (2)
In a UBGW tree whose reproduction law $\pi$, this translates into the following equation in law
[…]
where $N\sim\hat{\pi}$, the $w_{i}$’s and the $Z_{i}$’s are all independent, $w_{i}\sim\omega$ and $Z_{i}\sim Z$. As we will see, it is not very hard to show that this equation in law always has a solution, but the question of the uniqueness of the solution is trickier. In fact we were unable to prove it by conventional methods. We will come back to this in a few lines.
One of the main advances of the present work is that the matching $\mathbb{M}$ constructed from any solution of (4) by the decision rule is optimal. Our proof of this claim is inspired by Aldous’ work [aldous2000zeta2] on the Poisson Weighted Infinite Tree (PWIT). However, we work in the more general context of unimodular random trees, and one of our main difficulty is that, contrary to Aldous’ work, we deal with partial matchings rather than perfect matchings. Furthermore, there is no explicit solution to Equation (4). We overcome these two difficulties by introducing self-loops in Subsection 3.2 and by proving just enough properties on solutions to Equation (4) to allow us to conclude.
The second main contribution of Theorem 1, proved in Subsections 3.3 and 3.4, is that the law of $\mathbb{M}$ is unique and does not depend on the particular solution used. This bypasses previous restrictions on the belief propagation approach as in Gamarnik, Nowick and Swirscsz  [gamarnik2003maximum], Aldous and Steele [Aldous2004] or Aldous and Bandyopadhyay [Aldous_2005] that mostly required the variables $Z(u,v)$ to be measurable with respect to the tree.
The last step in proving Theorem 1 is to construct almost optimal matchings of finite graphs from an optimal matching $\mathbb{M}$ on their limiting tree. This is the purpose of section 4. A similar program has already been carried out by Aldous in the special case of the random assignment problem in bipartite graphs [Aldous1992AsymptoticsIT].
Let us come back to the uniqueness in law of the solution to Equation (4). As mentioned, we are not able to prove it directly and, quite surprisingly, we need the uniqueness of the optimal matching in Theorem 1 to establish it (under mild assumptions on the weights). This is formulated in the following theorem, proved in Section 5. Conversely, we emphasise that assuming uniqueness of the solution of Equation (4) does not help in proving Theorem 1.
Theorem 2.
@@
Let $\omega$ be a non atomic law such that there exists $a<\sup(\mathrm{supp}(\omega))$ such that $\left[a,\sup(\mathrm{supp}(\omega))\right)\subseteq\mathrm{supp}(\omega)$. Let $\hat{\pi}$ be a law on natural integers that has finite expectation. Then there is a unique law $\zeta$ satisfying:
@@
Corollary 3.
Let $G_{n}=(V_{n},E_{n},w_{n},o_{n})$ be a sequence of uniformly rooted weighted graphs as in Theorem 1. Take Z and Z’ to be independent copies of random variables satisfying (4) and $W$ with law $\omega$ independent of $Z$ and $Z^{\prime}$. One has
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\sum_{e\in\mathbb{M}_{\mathrm{opt}}(G_{n})}w(e)}{|E_{n}|}\right]$ $=\mathbb{E}[w(o)\mathbbm{1}_{o\in\mathbb{M}}]=\mathbb{E}\left[W\mathbbm{1}_{Z+Z^{\prime}<W}\right],$ (5)
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\left|E\left(\mathcal{M}_{\mathrm{opt}}(G_{n})\right)\right|}{|E_{n}|}\right]$ $=\mathbb{E}[\mathbbm{1}_{o\in\mathbb{M}}]=\mathbb{E}\left[\mathbbm{1}_{Z+Z^{\prime}<W}\right]=\frac{1-\phi(\hat{\phi}^{-1}(\mathbb{P}(Z=0))}{\phi^{\prime}(1)}.$ (6)
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\left|V\left(\mathcal{M}_{\mathrm{opt}}(G_{n})\right)\right|}{|V_{n}|}\right]=1-\phi\left(\hat{\phi}^{-1}(\mathbb{P}(Z=0))\right).$
We can also establish a conditional version of this statement:
Corollary 4.
Let $G_{n}=(V_{n},E_{n},w_{n},o_{n})$ be a sequence of uniformly rooted weighted graphs as in Theorem 1. Recall that $o_{n}$ is a uniform vertex of $G_{n}$. One has
The article is organised as follows. In Section 2, we introduce the notions of unimodular matchings and unimodular graphs. Section 3 carries out the study of optimal matchings on the limiting tree. In Section 4 we construct almost optimal matchings on finite graphs from their optimal infinite counterparts. In Section 5, we use the results established in the previous two sections to prove the two main theorems. Finally, Section 6 covers corollaries of Theorem 1 and a few adjacent problems to maximum matching on which this method also works.

### v2

@@ Abstract
We consider sequences of finite weighted random graphs that converge locally to unimodular i.i.d. weighted random trees. When the weights are atomless, we prove that the matchings of maximal weight converge locally to a matching on the limiting tree. For this purpose, we introduce and study unimodular matchings on weighted unimodular random trees as well as a notion of optimality for these objects. In this context, we prove that, in law, there is a unique optimal unimodular matching for a given unimodular tree. We then prove that this law is the local limit of the sequence of matchings of maximal weight. Along the way, we also show that this law is characterised by an equation derived from a message-passing algorithm.
@@ 1 Introduction
Optimal matchings in random graphs is a problem with a long history dating back to the paper by Karp and Sipser [4568355]. In this seminal paper, they obtain the asymptotic size of a maximal matching in a sparse Erdős-Rényi random graph with average degree $c\in(0,+\infty)$. Their method relies on the analysis of a greedy algorithm that constructs a matching in a graph. It turns out that the algorithm is simple when $c\leq e$ and less straightforward when $c>e$, making the analysis trickier in this regime.
@@
More recently, Gamarnik, Nowick and Swirscsz [gamarnik2003maximum] studied the edge-weighted version of this maximal matching problem on the Erdős-Rényi random graph and on regular random graphs. Their approach does not rely on the Karp–Sipser algorithm, but on a local optimality ansatz inspired by the earlier introduction of the so-called objective method by Aldous [Aldous2004, aldous2018processes, Aldous1992AsymptoticsIT, aldous2000zeta2]. Gamarnik, Nowick and Swirscsz were able to compute the asymptotic weight of the optimal matching in Erdős-Rényi random graph when $c\leq e$ and in the special case where the weights have exponential distribution.
The fundamental tool for studying these kinds of problems is the belief propagation algorithm. Informally, this algorithm assigns numbers to all edges of the graph, with an explicit dependency between neighbours, hence the name “propagation”. Moreover, they can be interpreted as beliefs: the lower the number, the more likely an edge belongs to the optimal matching. In homogeneous random graphs, the above dependency gives a fixed point equation in law for the beliefs, called the “belief propagation equation”. The restriction $c\leq e$ in [gamarnik2003maximum] comes from the fact that the employed method takes as input the uniqueness of solutions of iterates of the belief propagation equation.
One of the principles of Aldous’s objective method is that some asymptotic results are manifestations of limiting objects. In many settings involving random graphs, the objects are limits in law of the graphs for the local topology. In the special case where graphs are rooted uniformly, the graphs and their limits are called unimodular. This line of research has found tremendous success since the seminal paper by Benjamini and Schramm [BenjaminiSchramm1].
In this article, we study optimal matchings on weighted random graphs that converge to a random tree with iid weights in the Benjamin–Schramm sense. Note that we only require the weights to be independent in the limit. This includes classical models of random graphs such as Erdős-Rényi random graphs, regular random graphs or configuration models that all converge to unimodular versions of Bienaymé Galton Watson random trees. See for instance the books by van der Hofstad [vanderHofstad1, vanderHofstadBook2], Benjamini Lyons and Schramm [BENJAMINI_LYONS_SCHRAMM_2015], or Aldous and Lyons [aldous2018processes], and references therein.
@@
Strikingly, in this setting and under mild assumptions, the uniqueness of the solution to the iterates of the belief propagation equation mentioned above is not required for the convergence in law of the optimal matching. Informally, the limit $\mathbb{M}$ is a random matching on the unimodular limiting tree $\mathbb{T}$ rooted at a distinguished oriented edge $\overset{\rightarrow}{o}=(o_{-},o_{+})\in\overset{\rightarrow}{E}(\mathbb{T})$ and with i.i.d. edge-weights $w=(w(e))_{e\in E(\mathbb{T})}$ with law $\omega$, whose law is invariant by re-rooting. We call such random matchings unimodular (see Section 2 for precise definitions). In addition, the limit $\mathbb{M}$ is optimal in the sense that the quantity
@@
$\operatorname{perf}(\mathbb{T},\overset{\rightarrow}{o},\mathbb{M})$ $:=\mathbb{E}\left[w(\overset{\rightarrow}{o})\mathbbm{1}_{\overset{\rightarrow}{o}\in\mathbb{M}}\right]$ (1)
is maximal among all unimodular matchings of $(\mathbb{T},\overset{\rightarrow}{o},w)$. Note that for finite graphs, when $\overset{\rightarrow}{o}$ is chosen uniformly among directed edges, this performance is maximal when the total weight of the matching is maximal.
@@
Before stating our main result, we need to define more precisely Unimodular Bienaymé Galton Watson random trees (UBGW). Let $\pi$ be a distribution on non-negative integers with finite variance and generating function $\phi$. We denote by $\hat{\pi}$ the size-biased version of $\pi$, whose generating function is $\hat{\phi}$, defined by $\hat{\phi}(x)=\frac{\phi^{\prime}(x)}{\phi^{\prime}(1)}$. We say that a rooted tree $(\mathbb{T},\overset{\rightarrow}{o})$ is a UBGW with offspring distribution $\pi$ if the law of the tree can be obtained as follows. Take $(\mathbb{T}_{-},o_{-})$ and $(\mathbb{T}_{+},o_{+})$ two vertex-rooted independent Bienaymé-Galton-Watson trees with offspring distribution $\hat{\pi}$ (that is, each vertex of these trees have an independent number of children distributed according to $\hat{\pi}$) . The tree $(\mathbb{T},\overset{\rightarrow}{o})$ is then obtained by connecting $(\mathbb{T}_{-},o_{-})$ and $(\mathbb{T}_{+},o_{+})$ with the additional edge $\overset{\rightarrow}{o}=(o_{-},o_{+})$. We can now state our first main result:
Theorem 1 (Existence and uniqueness of the optimal matching for iid weighted UBGW).
@@
Let $(\mathbb{T},\overset{\rightarrow}{o},w)$ be an edge-rooted Unimodular Bienaymé-Galton-Watson (UBGW) tree with reproduction law $\pi$ and i.i.d edge-weights with law $\omega$ such that:
  • both $\pi$ and $\omega$ have finite expectation,
  • the law $\omega$ is atomless.
There exists a unique (in law) optimal matching $\mathbb{M}_{\mathrm{opt}}(\mathbb{T})$ on $(\mathbb{T},\overset{\rightarrow}{o},w)$. Furthermore, the distribution of $(\mathbb{T},\overset{\rightarrow}{o},w,\mathbb{M}_{\mathrm{opt}}(\mathbb{T}))$ is explicitly described in Proposition 3.2 (ii).
The description of $\mathbb{M}_{\mathrm{opt}}(\mathbb{T})$ given in Proposition 3.2 (ii) relies on the belief propagation algorithm mentioned above. In the simple case of a finite deterministic tree, the algorithm assigns to any oriented edge $(u,v)$ a belief variable $Z(u,v)$ that captures the penalty of excluding $v$ in the optimal matching of $T_{(u,v)}$, where $T_{(u,v)}$ is the connected component of $T\setminus\{\{u,v\}\}$ containing $v$. When $T_{(u,v)}$ is finite, this corresponds to the difference between the maximal weight of a matching of $T_{(u,v)}$ and the maximal weight of a matching of $T_{(u,v)}\setminus\{v\}$. From this, we can infer (see Section 3 for details) the so-called decision rule of the belief propagation procedure:
$\{u,v\}\in\mathbb{M}_{\mathrm{opt}}(\mathbb{T})\Leftrightarrow Z(u,v)+Z(v,u)<w(u,v).$ (2)
On an infinite random tree, such variables can still be defined, and this will be done in detail in Section 3.1. In a UBGW tree with reproduction law $\pi$ and i.i.d weights of law $\omega$, the stationary distribution satisfies the following equation in law:
[…]
where $N\sim\hat{\pi}$, the $w_{i}$’s and the $Z_{i}$’s are all independent, $w_{i}\sim\omega$ and $Z_{i}\sim Z$.
As we shall see, it is not very hard to show that this equation in law always has a solution, but the question of its uniqueness is trickier. It is not hard to see that any solution to Equation (4) gives with the decision rule of Equation (2) a unimodular matching of the underlying tree. Surprisingly, it turns out that the matching associated to any solution of Equation (4) is optimal.
Our second main result states the local convergence of maximal matchings on random graphs:
Theorem 2 (Local convergence of maximal matchings).
Let $G_{n}=(V_{n},E_{n},w_{n})$ be a sequence of weighted random graphs and let $\mathbb{M}_{\mathrm{opt}}(G_{n})$ be any maximum matching on $G_{n}$. Choose a root $\overset{\rightarrow}{o}_{n}$ uniformly among the directed edges and suppose that the sequence of rooted weighted random graphs $(G_{n},\overset{\rightarrow}{o}_{n},w_{n})$ converges locally to a vertex-rooted Unimodular Bienaymé-Galton-Watson (UBGW) tree $(\mathbb{T},\overset{\rightarrow}{o},w)$ with reproduction law $\pi$ and i.i.d edge-weights with law $\omega$ such that:
@@
Then $(G_{n},\overset{\rightarrow}{o}_{n},w_{n},\mathbb{M}_{\mathrm{opt}}(G_{n}))$ converges in law to $(\mathbb{T},\overset{\rightarrow}{o},w,\mathbb{M}_{\mathrm{opt}}(\mathbb{T}))$ for the local topology.
Corollary 1.
Let $G_{n}=(V_{n},E_{n},w_{n},\overset{\rightarrow}{o}_{n})$ be a sequence of uniformly rooted weighted graphs as in Theorem 2. Let $(\mathbb{T},w,\overset{\rightarrow}{o})$ be the local limit of the previous sequence. Take Z and Z’ to be independent copies of random variables satisfying (4) and $W$ with law $\omega$ independent of $Z$ and $Z^{\prime}$. One has
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\sum_{e\in E(\mathbb{M}_{\mathrm{opt}}(G_{n}))}w(e)}{|E_{n}|}\right]=\mathbb{E}\left[w(\overset{\rightarrow}{o})\mathbbm{1}_{\overset{\rightarrow}{o}\in\mathbb{M}_{\mathrm{opt}}(\mathbb{T})}\right]=\mathbb{E}\left[W\mathbbm{1}_{Z+Z^{\prime}<W}\right],$ (5)
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\left|E\left(\mathbb{M}_{\mathrm{opt}}(G_{n})\right)\right|}{|E_{n}|}\right]=\mathbb{P}\left(\overset{\rightarrow}{o}\in\mathbb{M}_{\mathrm{opt}(\mathbb{T})}\right)=\mathbb{P}\left(Z+Z^{\prime}<W\right)=\frac{1-\phi(\hat{\phi}^{-1}(\mathbb{P}(Z=0)))}{\phi^{\prime}(1)}.$ (6)
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\left|V\left(\mathbb{M}_{\mathrm{opt}}(G_{n})\right)\right|}{|V_{n}|}\right]=1-\phi\left(\hat{\phi}^{-1}(\mathbb{P}(Z=0))\right).$
We can also establish a conditional version of this statement:
Corollary 2.
Let $G_{n}=(V_{n},E_{n},w_{n})$ be a sequence of weighted graphs as in Theorem 2. Let $o_{n}$ be a uniform vertex of $G_{n}$. One has
With the aid of Theorem 1, we are able to give a probabilistic proof of the uniqueness of solutions to Equation (4):
Theorem 3 (Uniqueness of the belief propagation stationary measure).
@@
Let $\omega$ be a non atomic law that is not entirely singular with respect to Lebesgue measure in the neighbourhood of $\sup(\mathrm{supp}(\omega))$. Let $\hat{\pi}$ be a law on natural integers with finite expectation. Then there is a unique law $\zeta$ satisfying:
@@
As mentioned, we are not able to prove it directly, however, we are able to prove that for any matching distribution $\mathbb{M}$, there is at most one solution to Equation (4) that can yield $\mathbb{M}$ through the rule (2). Combined with the uniqueness of the optimal matching given by Theorem 1, we are then able to prove that there is only one solution to Equation (4). Conversely, we emphasise that assuming uniqueness of the solution of Equation (4) does not help in proving Theorem 1.
Let us conclude this introduction by mentioning a few papers that deal with matchings and Benjamini-Schramm convergence. Usually, the setting in these works consist of unweighted random graphs that converge for the topology and they prove the convergence of some matching statistics. Among these are the size of the maximal matchings [4568355, bordenave2012matchings] or the number of maximal matchings and the matching polynomial [matchingmeasure1, matchingmeasure2].
Organisation of the paper
In Section 2, we recall basic definitions on unimodularity and local convergence of random graphs. We then introduce the notions of unimodular matchings and optimality for unimodular matchings.
Section 3 carries out the study of optimal matchings on UBGW trees and contains the proof of Theorem 1. Our construction of unimodular matchings $\mathbb{M}$ is based on the message passing algorithm and inspired by Aldous’ work [aldous2000zeta2] on the Poisson Weighted Infinite Tree (PWIT). However, we work in the more general context of unimodular random trees, and one of the main difficulties is that, contrary to Aldous’ work, we deal with partial matchings rather than perfect matchings. Furthermore, there is no explicit solution to Equation (4). We overcome these two difficulties by introducing self-loops in Subsection 3.2 and by proving just enough properties on solutions to Equation (4). We then prove in Subsection 3.3 that the unimodular matching $\mathbb{M}$ we constructed is in fact optimal. Finally, in Subsection 3.4, we prove that the law of $\mathbb{M}$ is the only optimal matching law and does not depend on the particular solution of Equation (4) used. This bypasses previous restrictions on the belief propagation approach as in Gamarnik, Nowick and Swirscsz  [gamarnik2003maximum], Aldous and Steele [Aldous2004] or Aldous and Bandyopadhyay [Aldous_2005] that mostly required the variables $Z(u,v)$ to be measurable with respect to the tree.
The main idea behind Theorem 2 is to construct almost optimal matchings of finite graphs from an optimal matching $\mathbb{M}_{\mathrm{opt}}(\mathbb{T})$ on their limiting tree. This is the purpose of section 4. A similar program has already been carried out by Aldous in the special case of the random assignment problem on bipartite graphs [Aldous1992AsymptoticsIT].
In section 5, we deduce Theorem 3 by studying path conditionings on $(\mathbb{T},o,w,\mathbb{M}_{\mathrm{o}pt}(\mathbb{T}))$, in particular we show that if $\mathbb{M}_{\mathrm{opt}}(\mathbb{T})$ is given by a family of variables $(Z(u,v))_{(u,v)\in\overset{\rightarrow}{E}}$ through the rule (2), then the conditional law of the matching with respect to given weights on paths is sufficient to recover the law of the variables $(Z(u,v))_{(u,v)\in\overset{\rightarrow}{E}}$.
In section 6, we show how to apply Theorem 2 to obtain the announced corollaries. Finally, in Section 7, we discuss a few generalisations of the method to adjacent problems such as when the weights are on vertices instead of edges.